A basketball player sinks 50% of the 1500 shots he takes at the basket over a season. In the next season, he sinks 48% of the 1260 shots at basket. What is the probability that the performance in season 2 is better than that in season 1?
Answer #1

Result:

A basketball player sinks 50% of the 1500 shots he takes at the basket over a season. In the next season, he sinks 48% of the 1260 shots at basket. What is the probability that the performance in season 2 is better than that in season 1?

We have to find P( p2>p1) = P( p2-p1)>0

P= P( z=((p2-p1)/se >0) = P( z >-1.047)

P=0.8524

Hypothesis test for two independent proportions

P2

P1

pc

0.48

0.5

0.4909

p (as decimal)

605/1260

750/1500

1355/2760

p (as fraction)

604.8

750.

1354.8

X

1260

1500

2760

n

-0.02

Proportion difference

0.

hypothesized difference

0.0191

std. error

-1.047

z

.8524

p-value (one-tailed, upper)